OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2 A. x = (x – 6) + 2 B. x + y = y + 2 C. x + (y + 2) = 6 D. (y + 2) + y = 6<--

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Why do you think that? We know the value of x, so there should be no x's once we substitute.

OpenStudy (anonymous):

Oh so C

OpenStudy (sleepyjess):

There is an x in C :)

OpenStudy (usukidoll):

it just says Which equation shows the substitution method being used to solve the system of linear equations?... not solve for it??

OpenStudy (sleepyjess):

But yes, you are correct.

OpenStudy (anonymous):

:))) Thanx @sleepyjess

OpenStudy (usukidoll):

D is correct. for anyone curious... 2y+2=6 2y=4 y =2 x+2=6 x=4

OpenStudy (anonymous):

I'll tell you guys if your right

OpenStudy (anonymous):

But i 'm gonna choose C

OpenStudy (sleepyjess):

ugh I'm getting confused now, @UsukiDoll is correct

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

That was my answer at first tho -.-

OpenStudy (usukidoll):

seeeeee.............. re read the question Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2 If we let x = y+2, we have (y+2)+y=6

OpenStudy (anonymous):

Thank you @UsukiDoll

OpenStudy (usukidoll):

I miss doing easy stuff... I'm screwed with hard math D: not fun depending on the subject. Euclidean Geometry is fun.. Mathematical Bio... is a pain-o

OpenStudy (sleepyjess):

same, I'm in pre-calc right now, and I hate it